OpenStudy (anonymous):

The coast of the school banquet is $75+30n,where in is is the number of people of attending .What is the cost for 53 people

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

where you see \("n"\) replace it by \("53"\)

OpenStudy (anonymous):

ok what is the next thing i need healp fast

OpenStudy (jack1):

next u solve $75+30n $75+30(53) =?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

im kinda new here

OpenStudy (jack1):

all good @Shynice1234 ;)

OpenStudy (anonymous):

Im srry but i still dont get it im so srry ill leave

OpenStudy (jack1):

$75+30(53) = $75+30 x 53 = $75+1590 =?

OpenStudy (anonymous):

i got it now thanks Jack1

OpenStudy (jack1):

sweet as ;)

OpenStudy (anonymous):

How u make a queston